OpenStudy (anonymous):

Simplify 18 - 2[x + (x - 5)].

OpenStudy (anonymous):

for simplifying dis.u must remember BODMAS rule.

OpenStudy (anonymous):

18-2[x+(x-5)] =18-[2x+2(x-5)] =18-[2x+2x-10] =18-[4x-10] =18-4x+10 =18+10-4x =28-4x

OpenStudy (anonymous):

B=BRACKET O=OF D=DIVISION M=MULTIPLICATION A=ADDITION S=SUBTRACTION